​(Common stock valuation​) Abercrombie​ & Fitch's common stock pays a dividend of $1.50. It is currently selling for $36.15. If the​ firm's investors require a return of 11 percent on their investment from buying Abercrombie​ & Fitch​ stock, what growth rate would Abercrombie​ & Fitch have to provide the​ investors?
Answer #1

Current Dividend = D0 = $ 1.5, Required Return = R = 11% , Current Stock Price = P0 = $ 36.15 and let the required growth rate be G

Expected Dividend = D1 = D0 x (1+G) = 1.5 + 1.5G

Now, P0 = D1 / (R-G)

36.15 = (1.5 + 1.5G) / (0.11 - G)

36.15 x 0.11 - 36.15 G = 1.5 + 1.5 G

3.9765 - 36.15G = 1.5 + 1.5G

3.9765 - 1.5 = 36.15 + 1.5 G

37.65 G = 2.4765

G = 2.4765 / 37.65 = 0.065777 or 6.5777 % ~ 6.58 %